# Complex analytic solutions for the TQG model

### v1

@@ Abstract.
We present a condition under which the thermal quasi-geostrophic (TQG) model possesses a solution that is holomorphic in time with values in the Gevrey space of complex analytic functions. This can be seen as the complex extension of the work by Levermore and Oliver (1997) for the generalized Euler equation but applied to the TQG model.
@@ 1. Introduction
There are several 2-dimensional models for geophysical fluid flows where horizontal length scales are very large compared to their vertical scale. One such model is the Thermal Quasi-Geostrophic (TQG) model where there is a near balance between the Coriolis force, the hydrostatic pressure gradient force and the buoyancy gradient force. We refer to [2, 3, 12] for the historical account, the derivation and some simulations. In precise form, the TQG model posed on the spacetime spatially periodic cylinder $I\times\mathbb{T}^{2}$ where $I:=(0,T),T>0$, is a coupled system of equations governed by the evolution of the buoyancy $b:(t,\mathbf{x})\in I\times\mathbb{T}^{2}\mapsto b(t,\mathbf{x})\in\mathbb{R}$ and the potential vorticity $q:(t,\mathbf{x})\in I\times\mathbb{T}^{2}\mapsto q(t,\mathbf{x})\in\mathbb{R}$ in the following way:
@@
where $W^{3,2}_{\mathrm{div}_{\mathbf{x}}}(\mathbb{T}^{2})$ is the space of divergence-free vector fields in $W^{3,2}(\mathbb{T}^{2})$. Since we work on a torus, one can use the standard method of taking partial spatial derivative of the equation and applying commutator estimates to the advection terms (see [20, Proposition 2.1] and [13, Appendix] for instance) to extend this solution to $(b(t),q(t))\in W^{k+1,2}(\mathbb{T}^{2})\times W^{k,2}(\mathbb{T}^{2})$ for all $k\geq 2$. Consequently, smooth solutions $(b(t),q(t))\in C^{\infty}(\mathbb{T}^{2})\times C^{\infty}(\mathbb{T}^{2})$ also exists locally in time provided the data are also smooth. The existence of solutions in the larger class $(b(t),q(t))\in W^{k+1,2}(\mathbb{T}^{2})\times W^{k,2}(\mathbb{T}^{2})$ for any $k<2$ as well as the existence of global-in-time solution in any class, however, remains interesting open problems.
In this work, we are interested in studying the analytical property of smooth solutions of (1.1)-(1.4). In particular, we give a condition under which smooth solutions have convergent Taylor series. This will follow by showing that smooth solutions belongs to the Gevrey class of analytic functions under the condition that the radius of analyticity solves a suitable ordinary differential equation. The precise statement of this result is given below in Theorem 3.1. See [4] for a related result for the supercritical surface quasi-geostrophic equation, [14] for the hydrostatic Euler equations and [1, 5, 6, 15, 16, 17] for the (generalized) Euler equation
@@ 2.1. Notation
We consider a spacetime cylinder consisting of spatial points $\mathbf{x}=(x_{1},\ldots,x_{d})$ on the torus $\mathbb{T}^{d}=(\mathbb{R}/2\pi\mathbb{Z})^{d}$, $d=2,3$ with periodic boundary condition and a time variable $t\in[0,T]$ where $T>0$ is fixed and arbitrary. For functions $F$ and $G$, we write $F\lesssim G$ if there exists a generic constant $c>0$ such that $F\leq c\,G$. We also write $F\lesssim_{p}G$ if the constant $c(p)>0$ depends on a variable $p$. If $F\lesssim G$ and $G\lesssim F$ both hold (respectively, $F\lesssim_{p}G$ and $G\lesssim_{p}F$), we use the notation $F\sim G$ (respectively, $F\sim_{p}G$). The symbol ||$\cdot$|| ⋅ | may be used in four different context. For a scalar function $f\in\mathbb{R}$, $|f|$ denotes the absolute value of $f$. For a vector $\mathbf{f}\in\mathbb{R}^{d}$, $|\mathbf{f}|$ denotes the Euclidean norm of $\mathbf{f}$. For a square matrix $\mathbb{F}\in\mathbb{R}^{d\times d}$, $|\mathbb{F}|$ shall denote the Frobenius norm $\sqrt{\mathrm{trace}(\mathbb{F}^{T}\mathbb{F})}$. Also, if $S\subseteq\mathbb{R}^{d}$ is a (sub)set, then $|S|$ is the $d$-dimensional Lebesgue measure of $S$. Lastly, for any two normed spaces $(V_{1},\|\cdot\|_{V_{1}})$ and $(V_{2},\|\cdot\|_{V_{2}})$ with $v_{1}\in V_{1}$ and $v_{2}\in V_{2}$, we let
@@ 2.2. The functional setting
$\overline{\hat{\mathbf{f}}}_{\mathbf{k}}=\hat{\mathbf{f}}_{-\mathbf{k}},\quad\hat{\mathbf{f}}_{\bm{0}}=\bm{0},\quad\|\mathbf{f}\|_{H^{r}}=\bigg{(}\sum_{\mathbf{k}\in\mathbb{Z}^{d}_{\bm{0}}}|\mathbf{k}|^{2r}|\hat{\mathbf{f}}_{\mathbf{k}}|^{2}\bigg{)}^{\frac{1}{2}}<\infty\Bigg{\}}.$
Here, the $\hat{\mathbf{f}}_{\mathbf{k}}$s are the Fourier coefficients of the function $\mathbf{f}\in L^{2}(\mathbb{T}^{d})$ and $\mathbb{Z}^{d}_{\bm{0}}:=2\pi\mathbb{Z}^{d}\setminus\{\bm{0}\}$. Furthermore, note the $\hat{\mathbf{f}}_{\bm{0}}=\bm{0}$ is equivalent to saying that elements $\mathbf{f}\in H^{r}(\mathbb{T}^{d})$ in (2.1) are mean-free, i.e., $\int_{\mathbb{T}^{d}}\mathbf{f}(\mathbf{x})\,\mathrm{d}\mathbf{x}=0$.
@@
Here, the condition $\hat{\mathbf{f}}_{\mathbf{k}}\cdot\mathbf{k}=0$ corresponds to the divergence-free condition $\mathrm{div}_{\mathbf{x}}\mathbf{f}=0$ in frequency space. Thus, $\mathbb{H}^{r}(\mathbb{T}^{d})$ is the subclass of divergence-free vectors in $H^{r}(\mathbb{T}^{d})$. Next, if we consider the positive, self-adjoint operator $\Lambda=-\Delta$ which is known to posses a nondecreasing sequence $\{\lambda_{j}\}_{j\in\mathbb{N}}$ of strictly positive eigenvalues that approaches infinity as $j\rightarrow\infty$, see [18]. Then for any $r\in\mathbb{R}$, we can define its fractional power $\Lambda^{r}$ as the mapping
@@
$D(e^{\varphi\Lambda^{1/2s}}:\mathbb{H}^{r}(\mathbb{T}^{d}))\equiv\big{\{}\mathbf{f}(\mathbf{x})\in\mathbb{H}^{r}(\mathbb{T}^{d})\,:\,\|e^{\varphi\Lambda^{1/2s}}\mathbf{f}(\mathbf{x})\|_{H^{r}}<\infty\big{\}}$ (2.3)
@@
Lemma 2.1 generalises [17, Lemma 8] in two ways and we defer its proof to the appendix below. Firstly, nowhere is the incompressibility condition used (hence $\mathbf{u}\in D(e^{\varphi\Lambda^{1/2}}:{H}^{r+1/2}(\mathbb{T}^{d}))$ rather than $\mathbf{u}\in D(e^{\varphi\Lambda^{1/2}}:\mathbb{H}^{r+1/2}(\mathbb{T}^{d}))$) whereas $\mathbf{u}$ was taken to be (weighted) incompressible in the cited result. Secondly, we clearly do not require $\mathbf{v}=\mathbf{w}$ nor do they need to be scaler-valued.
Finally, we recall the following result in [8, Lemma 2.2]. Although the original result is stated for non-negative integers $k\in\mathbb{N}\cup\{0\}$, a look at the proof shows that it extends to all non-negative real numbers $r\geq 0$ with no further modication to the proof.
@@ 3.1. Complexified TQG
$\Theta(s):=$ $\|(e^{\varphi(s\cos\theta)\Lambda^{1/2}}b(\zeta,\cdot)\,,\,e^{\varphi(s\cos\theta)\Lambda^{1/2}}q(\zeta,\cdot))\|_{H^{4}_{\mathbb{C}}\times H^{3}_{\mathbb{C}}}$ (3.8)
@@
Under the Beale–Kato–Majda criterion for the blow-up of solution to the TQG [9], the time span $T>0$ can be made arbitrary large. In this case, the complex neighbourhood where the solution is of Gevrey analytic class becomes
$\left\{\zeta=se^{i\theta}\quad:\quad\mathrm{Re}\,\zeta>0,\quad|\theta|<\frac{\pi}{2},\quad 0<s\,\mathcal{D}(\mathrm{data})^{\frac{1}{2}}<\frac{1}{c\cos\theta}\right\}$
so that by tuning $\theta$, one can make $s$ arbitrary large.
@@ 3.2. Proof of result
We now present the proof of Theorem 3.1 which uses arguments in [11, 17] and [10, Chapter II]. This will involve formal a priori estimates which, as usual, can be made rigorous by first considering a finite-dimensional Galerkin approximation and later passing to the limit in the approximation parameter. With this said, for $\varphi:=\varphi(\mathrm{Re}\,\zeta)=\varphi(s\cos\theta)$ with fixed $\theta\in(-\frac{\pi}{2},\frac{\pi}{2})$, we first consider the following equation that is due to the product rule for the product of two functions:
@@
$\quad+\|e^{\varphi\Lambda^{1/2}}\mathbf{u}\|_{H^{9/2}_{\mathbb{C}}}\|e^{\varphi\Lambda^{1/2}}b\|_{H^{4}_{\mathbb{C}}}\|e^{\varphi\Lambda^{1/2}}b\|_{H^{9/2}_{\mathbb{C}}}$
@@
$\quad+\|e^{\varphi\Lambda^{1/2}}\mathbf{u}\|_{H^{7/2}_{\mathbb{C}}}\|e^{\varphi\Lambda^{1/2}}q\|_{H^{3}_{\mathbb{C}}}\|e^{\varphi\Lambda^{1/2}}q\|_{H^{7/2}_{\mathbb{C}}}$
@@ 4. Appendix
We devote this section to the proof of Lemma 2.1 which follows the arguments of [17, Lemma 8]. We rely on two preliminary results, the first of which is given by [17, Lemma 9].
@@
where in the last line, we have used Young inequality to obtain
@@
as well as the identity $\hat{\mathbf{z}}_{\bm{\ell}}=\hat{\mathbf{z}}_{-(\mathbf{j}+\mathbf{k})}$ which hold since $\mathbf{j}+\mathbf{k}+\bm{\ell}=\bm{0}$. Consequently, we obtain

### v2

@@ Abstract.
We present a condition under which the thermal quasi-geostrophic (TQG) model possesses a solution that is holomorphic in time with values in the Gevrey space of complex analytic functions. This can be seen as the complex extension of the work by Levermore and Oliver (1997) for the generalised Euler equation but applied to the TQG model.
@@ 1. Introduction
There are several 2-dimensional models for geophysical fluid flows where horizontal length scales are very large compared to their vertical scale. One such model is the Thermal Quasi-Geostrophic (TQG) model where there is a near balance between the Coriolis force, the hydrostatic pressure gradient force and the buoyancy gradient force. We refer to [2, 3, 13] for the historical account, the derivation and some simulations. In precise form, the TQG model posed on the space-time spatially periodic cylinder $I\times\mathbb{T}^{2}$ where $I:=(0,T),T>0$, is a coupled system of equations governed by the evolution of the buoyancy $b:(t,\mathbf{x})\in I\times\mathbb{T}^{2}\mapsto b(t,\mathbf{x})\in\mathbb{R}$ and the potential vorticity $q:(t,\mathbf{x})\in I\times\mathbb{T}^{2}\mapsto q(t,\mathbf{x})\in\mathbb{R}$ in the following way:
@@
where $W^{3,2}_{\mathrm{div}_{\mathbf{x}}}(\mathbb{T}^{2})$ is the space of divergence-free vector fields in $W^{3,2}(\mathbb{T}^{2})$. Since we work on a torus, one can use the standard method of taking partial spatial derivative of the equation and applying commutator estimates to the advection terms (see [21, Proposition 2.1] and [14, Appendix] for instance) to extend this solution to
$(b(t),q(t))\in W^{k+1,2}(\mathbb{T}^{2})\times W^{k,2}(\mathbb{T}^{2})\quad\text{for all }k\geq 2.$
Consequently, smooth solutions $(b(t),q(t))\in C^{\infty}(\mathbb{T}^{2})\times C^{\infty}(\mathbb{T}^{2})$ also exist locally in time provided the data are also smooth. The existence of solutions in the larger class $(b(t),q(t))\in W^{k+1,2}(\mathbb{T}^{2})\times W^{k,2}(\mathbb{T}^{2})$ for any $k<2$ as well as the existence of global-in-time solution in any class, however, remain interesting open problems.
In this work, we are interested in studying the analytical property of smooth solutions of (1.1)-(1.4). In particular, we give a condition under which smooth solutions have convergent Taylor series. This will follow by showing that smooth solutions belong to the Gevrey class of analytic functions under the condition that the radius of analyticity solves a suitable ordinary differential equation. The precise statement of this result is given below in Theorem 3.1. See [4] for a related result for the supercritical surface quasi-geostrophic equation, [15] for the hydrostatic Euler equations and [1, 5, 6, 16, 17, 18] for the (generalised) Euler equation
@@ 2.1. Notation
We consider a space-time cylinder consisting of spatial points $\mathbf{x}=(x_{1},\ldots,x_{d})$ on the torus $\mathbb{T}^{d}=(\mathbb{R}/2\pi\mathbb{Z})^{d}$, $d=2,3$ with periodic boundary condition and a time variable $t\in[0,T]$ where $T>0$ is fixed and arbitrary. For functions $F$ and $G$, we write $F\lesssim G$ if there exists a generic constant $c>0$ such that $F\leq c\,G$. We also write $F\lesssim_{p}G$ if the constant $c(p)>0$ depends on a variable $p$. If $F\lesssim G$ and $G\lesssim F$ both hold (respectively, $F\lesssim_{p}G$ and $G\lesssim_{p}F$), we use the notation $F\sim G$ (respectively, $F\sim_{p}G$). The symbol ||$\cdot$|| ⋅ | may be used in four different contexts. For a scalar function $f\in\mathbb{R}$, $|f|$ denotes the absolute value of $f$. For a vector $\mathbf{f}\in\mathbb{R}^{d}$, $|\mathbf{f}|$ denotes the Euclidean norm of $\mathbf{f}$. For a square matrix $\mathbb{F}\in\mathbb{R}^{d\times d}$, $|\mathbb{F}|$ shall denote the Frobenius norm $\sqrt{\mathrm{trace}(\mathbb{F}^{T}\mathbb{F})}$. Also, if $S\subseteq\mathbb{R}^{d}$ is a (sub)set, then $|S|$ is the $d$-dimensional Lebesgue measure of $S$. Lastly, for any two normed spaces $(V_{1},\|\cdot\|_{V_{1}})$ and $(V_{2},\|\cdot\|_{V_{2}})$ with $v_{1}\in V_{1}$ and $v_{2}\in V_{2}$, we let
@@ 2.2. The functional setting
$\overline{\hat{\mathbf{f}}}_{\mathbf{k}}=\hat{\mathbf{f}}_{-\mathbf{k}},\quad\hat{\mathbf{f}}_{\bm{0}}=\bm{0},\quad\|\mathbf{f}\|_{H^{r}}=\bigg{(}\sum_{\mathbf{k}\in\mathbb{Z}^{d}_{\bm{0}}}|\mathbf{k}|^{2r}|\hat{\mathbf{f}}_{\mathbf{k}}|^{2}\bigg{)}^{1/2}<\infty\Bigg{\}}.$
Here, the $\hat{\mathbf{f}}_{\mathbf{k}}$s are the Fourier coefficients of the function $\mathbf{f}\in L^{2}(\mathbb{T}^{d})$ and $\mathbb{Z}^{d}_{\bm{0}}:=2\pi\mathbb{Z}^{d}\setminus\{\bm{0}\}$. Furthermore, note that $\hat{\mathbf{f}}_{\bm{0}}=\bm{0}$ is equivalent to saying that elements $\mathbf{f}\in H^{r}(\mathbb{T}^{d})$ in (2.1) are mean-free, i.e., $\int_{\mathbb{T}^{d}}\mathbf{f}(\mathbf{x})\,\mathrm{d}\mathbf{x}=0$.
@@
Here, the condition $\hat{\mathbf{f}}_{\mathbf{k}}\cdot\mathbf{k}=0$ corresponds to the divergence-free condition $\mathrm{div}_{\mathbf{x}}\mathbf{f}=0$ in frequency space. Thus, $\mathbb{H}^{r}(\mathbb{T}^{d})$ is the subclass of divergence-free vectors in $H^{r}(\mathbb{T}^{d})$. Next, if we consider the positive, self-adjoint operator $\Lambda=-\Delta$ which is known to possess a nondecreasing sequence $\{\lambda_{j}\}_{j\in\mathbb{N}}$ of strictly positive eigenvalues that approach infinity as $j\rightarrow\infty$, see [19]. Then for any $r\in\mathbb{R}$, we can define its fractional power $\Lambda^{r}$ as the mapping
@@
$D(e^{\varphi\Lambda^{1/2s}}:\mathbb{H}^{r}(\mathbb{T}^{d}))\equiv\big{\{}\mathbf{f}(\mathbf{x})\in\mathbb{H}^{r}(\mathbb{T}^{d})\,:\,\|e^{\varphi\Lambda^{1/2s}}\mathbf{f}(\mathbf{x})\|_{H^{r}}<\infty\big{\}}$
@@
Lemma 2.1 generalises [18, Lemma 8] in two ways and we defer its proof to the appendix below. Firstly, nowhere is the incompressibility condition used (hence $\mathbf{u}\in D(e^{\varphi\Lambda^{1/2}}:{H}^{r+1/2}(\mathbb{T}^{d}))$ rather than $\mathbf{u}\in D(e^{\varphi\Lambda^{1/2}}:\mathbb{H}^{r+1/2}(\mathbb{T}^{d}))$) whereas $\mathbf{u}$ was taken to be (weighted) incompressible in the cited result. Secondly, we clearly do not require $\mathbf{v}=\mathbf{w}$ nor do they need to be scaler-valued.
Finally, we recall the following result in [8, Lemma 2.2]. Although the original result is stated for non-negative integers $k\in\mathbb{N}\cup\{0\}$, a look at the proof shows that it extends to all non-negative real numbers $r\geq 0$ with no further modification to the proof.
@@ 3.1. Complexified TQG
$\Theta(s):=$ $\|(e^{\varphi(s\cos\theta)\Lambda^{1/2}}b(\zeta,\cdot)\,,\,e^{\varphi(s\cos\theta)\Lambda^{1/2}}q(\zeta,\cdot))\|_{H^{4}_{\mathbb{C}}\times H^{3}_{\mathbb{C}}}$
@@
Under the Beale–Kato–Majda criterion for the blow-up of solution to the TQG [9], the time span $T>0$ can be made arbitrarily large. In this case, the complex neighbourhood where the solution is of Gevrey analytic class becomes
$\left\{\zeta=se^{i\theta}\quad:\quad\mathrm{Re}\,\zeta>0,\quad|\theta|<\frac{\pi}{2},\quad 0<s\,\mathcal{D}(\mathrm{data})^{1/2}<\frac{1}{c\cos\theta}\right\}$
so that by tuning $\theta$, one can make $s$ arbitrarily large.
@@ 3.2. Proof of result
We now present the proof of Theorem 3.1 which uses arguments in [10, 12, 18] and [11, Chapter II]. This will involve formal a priori estimates which, as usual, can be made rigorous by first considering a finite-dimensional Galerkin approximation and later passing to the limit in the approximation parameter. With this said, for $\varphi:=\varphi(\mathrm{Re}\,\zeta)=\varphi(s\cos\theta)$ with fixed $\theta\in(-\frac{\pi}{2},\frac{\pi}{2})$, we first consider the following equation that is due to the product rule for the product of two functions:
@@
$\quad+\|e^{\varphi\Lambda^{1/2}}\mathbf{u}\|_{H^{9/2}_{\mathbb{C}}}\|e^{\varphi\Lambda^{1/2}}b\|_{H^{4}_{\mathbb{C}}}\|e^{\varphi\Lambda^{1/2}}b\|_{H^{9/2}_{\mathbb{C}}}\big{)}$
@@
$\quad+\|e^{\varphi\Lambda^{1/2}}\mathbf{u}\|_{H^{7/2}_{\mathbb{C}}}\|e^{\varphi\Lambda^{1/2}}q\|_{H^{3}_{\mathbb{C}}}\|e^{\varphi\Lambda^{1/2}}q\|_{H^{7/2}_{\mathbb{C}}}\big{)}$
@@ 4. Appendix
We devote this section to the proof of Lemma 2.1 which follows the arguments of [18, Lemma 8]. We rely on two preliminary results, the first of which is given by [18, Lemma 9].
@@
where in the last line, we have used Young’s inequality to obtain
@@
as well as the identity $\hat{\mathbf{z}}_{\bm{\ell}}=\hat{\mathbf{z}}_{-(\mathbf{j}+\mathbf{k})}$ which holds since $\mathbf{j}+\mathbf{k}+\bm{\ell}=\bm{0}$. Consequently, we obtain